I think I have the equation incorrect. Please help. A dairy has 400 quarts of whole milk containing 5% butterfat. How many quarts of low-fat milk containing 1.5% butterfat should be added tp produce 2% butterfat.
I have 5X+1(400-X)=400(2)

Answer by Paul(988)

400(5)-1.5x=2(x+500)
2000-1.5x=2x+1000
-3.5x=-1000
x=285.7(rounded)

400-285.7=114.3

Hence, for 1.5% there is 285.7 Quarts and for 2% there is 114.3 quarts.
Paul.